# Structural complexity of snapshots of 2D Fermi-Hubbard systems

### v1

@@ Abstract
The development of quantum gas microscopy for two-dimensional optical lattices has provided an unparalleled tool to study the Fermi-Hubbard model (FHM) with ultracold atoms. Spin-resolved projective measurements, or snapshots, have played a significant role in quantifying correlation functions which uncover underlying physical phenomena such as antiferromagnetism at commensurate filling on bipartite lattices, and other charge and spin correlations, as well as dynamical properties at various densities. However, in order to interpret these results and to establish temperature scales, comparison against theory is generally needed. Here, we employ a recent concept, the multi-scale structural complexity, and show that when computed for the snapshots (of either single spin species, local moments, or total density) it can provide a theory-free property, immediately accessible to experiments. Specifically, after benchmarking results for Ising and XY models, we study the structural complexity of snapshots of the repulsive FHM in the two-dimensional square lattice as a function of doping and temperature. We generate projective measurements using determinant quantum Monte Carlo and compare their complexities against those from the experiment. We demonstrate that these complexities are linked to relevant physical observables such as the entropy and double occupancy. Their behaviors capture the development of correlations and relevant length scales in the system. We provide an open-source code in Python which can be implemented into data analysis routines in experimental settings for the square lattice [1].
@@ I Introduction
One of the major accomplishments in quantum simulation has been the precise engineering of the FHM with ultracold atoms in optical lattices. These exhibit flexible tunability of the kinetic and interaction energies, as well as the lattice filling or even the lattice geometry [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. In lower dimensional lattices, i.e. one-dimensional (1D) and 2D, quantum gas microscopy has provided direct observation of correlations beyond nearest neighbors through real-space imaging of spin-resolved projective measurements [19, 20, 21, 22, 23, 24, 25, 26, 27]. These results yield profound insights into the phase diagram [28], in particular regarding antiferromagnetism, and how spin correlations are affected by the presence of holes at intermediate temperatures.
@@ II.2 The Fermi-Hubbard model and the DQMC
$H=-t\sum_{\langle i,j\rangle,\sigma}\left(c_{i\sigma}^{\dagger}c_{j\sigma}+\mathrm{h.c.}\right)+U\sum_{i}n_{i\uparrow}n_{i\downarrow},-\mu\sum_{i,\sigma}n_{i\sigma},$ (2)
where $c_{i\sigma}^{\dagger}$ ($c_{i\sigma}$) is the creation (annihilation) operator for a fermion with spin flavor $\sigma=\uparrow,\downarrow$ on site $i=1,2,...,N$ in a 2D square lattice, $N$ denotes the number of lattice sites, $n_{i\sigma}=c_{i\sigma}^{\dagger}c_{i\sigma}$ is the number operator for flavor $\sigma$, $t$ is the nearest-neighbor hopping amplitude, and $U$ is the interaction strength. We work in the grand canonical ensemble and use the chemical potential $\mu$ to adjust the fermion density. We set $\hbar=k_{B}=1$ throughout the paper, and consider the $U>0$ (repulsive) case.
In this work, we generate snapshots of projective measurements in $10\times 10$ lattices using the method described in Ref. [44], in which nested componentwise direct sampling of fermion pseudodensity matrices is used to generate projective measurements of the occupation numbers. The procedure introduces a weight associated with each snapshot, which for all the cases we consider here is more or less uniform among snapshots, and we therefore ignore for simplicity. We use DQMC with a Trotter step of $\Delta\tau=0.05t$. The projective measurements correspond to the spin-resolved densities $n_{i\sigma}$ in the lattice, from which the local moment $m^{z}_{i}=|n_{i\uparrow}-n_{i\downarrow}|$ and the total density $n_{i}=n_{i\uparrow}+n_{i\downarrow}$ are constructed. We apply point-group symmetries to increase the number of samples eightfold.
@@ II.3 Structural complexity
is the overlap function, where the image has $N$ pixels, and $u^{(k)}_{i}$ corresponds to the value of the pixel at site $i$ at coarse-graining step $k$. The coarse-graining procedure using a $\Lambda\times\Lambda$ window and the calculation of the overlaps is depicted in Fig. 1. $D_{k}$ is the “dissimilarity” at coarse-graining step $k$. It contains correlations that extend up to a linear size of $\Lambda^{k}$ in the original image and measures how different the images at consecutive coarse-graining steps $k$ and $k+1$ are. Here, we set $\Lambda=2$. As we will see later, it is also useful to define $C_{1}=\mathcal{C}_{0}-D_{0}$ [i.e. we start the sum in Eq. (3) at $k=1$ rather than $k=0$].
@@ III Results
Figure 2: Classical models in $128\times 128$ lattices. Results are presented for Ising FM (a)-(b) and AFM (c)-(d) couplings, and for the XY model (e)-(f). Upper row: $\mathcal{C}_{0}$ (blue circles) and $\mathcal{C}_{1}$ (red squares) as a function of $T/J$. Solid lines correspond to the absolute value of the derivative. The extracted transition temperatures $T_{c}$ are consistent with the exact results in all cases. In panel (e) black markers and lines correspond tor results obtained in $256\times 256$ lattices. Lower row: $D_{k}$ vs $k$ for different temperatures. The behavior of the dissimilarities for each model are significantly different.
@@ III.1.2 The Hamiltonian with $O(2)$ symmetry
We now turn our attention to the XY model. In Fig. 2(e) we show the structural complexity of the model as a function of temperature for two system sizes $128\times 128$ (solid markers) and $256\times 256$ (open black markers). The results are averaged over five different initial random seeds. We further apply a moving average with a five-point window over temperature fitted with a local third-order polynomial (the Savitzky-Golay filter) to reduce noise. Similarly to the AFM Ising model, $\mathcal{C}_{0}$ is more or less constant for all temperatures and for both system sizes, indicating the near zero net magnetization of the system. On the contrary, $\mathcal{C}_{1}$ increases as the temperature is lowered, indicating that the low-$T$ phase, where the existence of pairs of vortices with opposite vorticity is expected [45], is the most complex one. We find a strong finite-size dependence in $\mathcal{C}_{1}$, expected to persist in this model for certain properties even with systems an order of magnitude larger in size [47]. The derivative of $\mathcal{C}_{1}$ for the $128\times 128$ system exhibits broad and inconclusive features at $T/J\lesssim 3$. By increasing the system size to $256\times 256$, these features develop into a sharp upturn at the lowest temperatures we could access in our simulations due to inflating autocorrelation times. This feature is consistent with the location of the peak in the specific heat ($T/J\sim 1.04$) and the BKT transition temperature of the model ($T_{c}/J=0.894$[48, 47].
@@ III.2.1 Spin-resolved densities $n_{\sigma}$
More interestingly, we find that $\mathcal{C}_{1}$ closely tracks the temperature and doping behavior of the entropy per site $S/N$ in the model (see Fig. 7). We observe that below $T/t<1$, the structural complexity develops a peak around 20% doping. Although the peak in the entropy occurs at a slightly lower doping (see the inset of Fig. 7), the shape and trends are fairly similar. Additionally, we also observe good agreement with the experimental data at all temperatures considered, which further supports the validity of our results. Because of the correlation between $\mathcal{C}_{1}$ and $S/N$, this suggests the structural complexity of Fermi-Hubbard snapshots can be used as a quick and rapid proxy for the entropy of the system. This can be very useful in optical lattice experiments, where after loading the lattice, the calculation of the entropy is not straightforward [50].
Another important observation is the fact that at the temperatures considered, the region where the structural complexity is maximal corresponds to the strange metal region. The broad peak in $\mathcal{C}_{1}$ spanning $\delta\in[10\%,30\%]$ is consistent with the doping value $\delta_{c}\sim 30\%$ that corresponds to the crossover to the Fermi liquid phase [51]. Moreover, the trends in the structural complexity vs doping illustrates that features related to phases that emerge in this doping range at lower temperatures (such as fluctuating stripes [29, 31, 30]) are more complex than the AFM parent state at half-filling, and that this metallic region is also more complex than the conventional Fermi liquid found at larger dopings.
@@ III.2.2 Local moments and total density
Before quantum gas microscopy had access to images of both fermionic species of the same system at the same time, density snapshots of the FHM could not be obtained; only local moments could be detected separately. Nowadays, with spin-resolved imaging, local moment as well as total density images can be created [52]. In this section we study their structural complexities and dissimilarities.
@@
In Fig. 9, we show the temperature dependence of the dissimilarities for the density snapshots at $\delta=0.18$. Here we observe that (i) $D_{k}$ decays rapidly with increasing $k$, (ii) $D_{0}$ exhibits a non monotonic behavior with $T/t$, akin to the double occupancy, $\mathcal{D}=\langle n_{i\uparrow}n_{i\downarrow}\rangle$, and (iii) it closely follows the function $\mathcal{D}^{\prime}=(3/4)[\mathcal{D}+0.5\delta(1-\delta)]$. This expression can be derived analytically under the assumption that the uncorrelated parts of the correlation functions dominate (see Appendix B), and it holds for other dopings too. The agreement at high-$T$, where the assumption is valid, illustrates that $D_{0}$ for the density snapshots is in fact closely related to the double occupancy fraction in the model. More important are the discrepancies at $T/t\leq 1$, which signal the presence of non-trivial 2-point nearest- and next-nearest-neighbor spin-resolved density-density correlations in the model. A similar approximation for $D_{0}$ based on local moment snapshots proves challenging as 3-point and 4-point density correlators enter the analysis (see Appendix B).
@@
Our results point to the existence of a possible scale invariance in the model at $\delta\sim 18\%$. Fig. 11 presents $2^{\langle k\rangle}$ vs doping at different temperatures for the three different types of snapshots analyzed. At the highest temperature considered, $T/t=4$, $2^{\langle k\rangle}$ increases slowly with $\delta$ for the $n_{\sigma}$ and $n$ snapshots, but decreases for the $m_{z}$ snapshots. This trend persists down to all temperatures considered for the $m_{z}$ snapshots, becoming sharper as the temperature is lowered, but surprisingly exhibiting an almost temperature-independent crossing at $\delta\sim 18\%$. This coincides with the region where structural complexity is maximal and corresponds to the strange metal region. Since this occurs at easily experimentally-accessible regimes, it is a phenomenon that can be explored immediately in the experiment.
$2^{\langle k\rangle}$ vs temperature for $n_{\sigma}$ and $n$ snapshots displays features at doings that are consistent with the above observations. As seen in Fig. 11(a) for $n_{\sigma}$, the curves develop a plateau that starts around $\delta\sim 18\%$ and extends up to $\delta\sim 40\%$, while for $n$ in Fig. 11(b) the curves develop a broad peak around $\delta\sim 18\%$.
@@ Acknowledgements.
We are grateful to Waseem Bakr, Elmer Guardado-Sanchez, and Benjamin M. Spar for sharing their experimental snapshots for the Fermi-Hubbard model with us. We also thank Waseem Bakr for his insightful comments on our manuscript. The authors are supported by the grant DE-SC-0022311, funded by the U.S. Department of Energy, Office of Science. EIGP would like to acknowledge useful discussions with Steven Johnston and Bejamin Cohen-Stead while visiting the University of Tennessee, Knoxville.

### v2

@@ Abstract
The development of quantum gas microscopy for two-dimensional optical lattices has provided an unparalleled tool to study the Fermi-Hubbard model (FHM) with ultracold atoms. Spin-resolved projective measurements, or snapshots, have played a significant role in quantifying correlation functions, theory verification, and thus, the uncovering of underlying physical phenomena such as antiferromagnetism at commensurate filling on bipartite lattices, and other charge and spin correlations, as well as dynamical properties at various densities. Here, we employ a recent concept, the multi-scale structural complexity, and show that when computed for the snapshots (of either single spin species, local moments, or total density) it can provide a theory-free property, immediately accessible to experiments. Specifically, after benchmarking results for Ising and XY models, we study the structural complexity of snapshots of the repulsive FHM in the two-dimensional square lattice as a function of doping and temperature. We generate projective measurements using determinant quantum Monte Carlo and compare their complexities against those from the experiment. We demonstrate that these complexities are linked to relevant physical observables such as the entropy and double occupancy. Their behaviors capture the development of correlations and relevant length scales in the system. We provide an open-source code in Python which can be implemented into data analysis routines in experimental settings for the square lattice [1].
@@ I Introduction
One of the major accomplishments in quantum simulation has been the precise engineering of the FHM with ultracold atoms in optical lattices. These exhibit flexible tunability of the kinetic and interaction energies, as well as the lattice filling or even the lattice geometry [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. In lower dimensional lattices, i.e. one-dimensional (1D) and 2D, quantum gas microscopy has provided direct observation of correlations beyond nearest neighbors through real-space imaging of spin-resolved projective measurements [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. These results yield profound insights into the phase diagram [30], in particular regarding antiferromagnetism, and how spin correlations are affected by the presence of holes at intermediate temperatures.
@@ II.2 The Fermi-Hubbard model and the DQMC
$H=-t\sum_{\langle i,j\rangle,\sigma}\left(c_{i\sigma}^{\dagger}c_{j\sigma}^{\phantom{\dagger}}+\mathrm{h.c.}\right)+U\sum_{i}n_{i\uparrow}n_{i\downarrow},-\mu\sum_{i,\sigma}n_{i\sigma},$ (2)
where $c_{i\sigma}^{\dagger}$ ($c_{i\sigma}^{\phantom{\dagger}}$) is the creation (annihilation) operator for a fermion with spin flavor $\sigma=\uparrow,\downarrow$ on site $i=1,2,...,N$ in a 2D square lattice, $N$ denotes the number of lattice sites, $n_{i\sigma}=c_{i\sigma}^{\dagger}c_{i\sigma}^{\phantom{\dagger}}$ is the number operator for flavor $\sigma$, $t$ is the nearest-neighbor hopping amplitude, and $U$ is the interaction strength. We work in the grand canonical ensemble and use the chemical potential $\mu$ to adjust the fermion density. We set $\hbar=k_{B}=1$ throughout the paper, and consider the $U>0$ (repulsive) case.
In this work, we generate snapshots of projective measurements in $10\times 10$ lattices using the method described in Ref. [46, 49], in which nested componentwise direct sampling of fermion pseudodensity matrices is used to generate an ensemble of pseudosnapshots, which, when reweighted, is equivalent to an ensemble of projective measurements of the occupation numbers. The procedure introduces a weight associated with each snapshot, which for all the cases we consider here is more or less uniform among snapshots, and we therefore ignore for simplicity. We use DQMC with a Trotter step of $\Delta\tau=0.05t$. The projective measurements correspond to the spin-resolved densities $n_{i\sigma}$ in the lattice, from which the local moment $m^{z}_{i}=|n_{i\uparrow}-n_{i\downarrow}|$ and the total density $n_{i}=n_{i\uparrow}+n_{i\downarrow}$ are constructed. We apply point-group symmetries to increase the number of samples eightfold.
@@ II.3 Structural complexity
is the overlap function, where the image has $N$ pixels, and $u^{(k)}_{i}$ corresponds to the value of the pixel at site $i$ at coarse-graining step $k$. The coarse-graining procedure using a $\Lambda\times\Lambda$ window and the calculation of the overlaps is depicted in Fig. 1. $D_{k}$ is the “dissimilarity” at coarse-graining step $k$. It contains correlations that extend up to a linear size of $\Lambda^{k}$ in the original image and measures how different the images at consecutive coarse-graining steps $k$ and $k+1$ are. Here, we set $\Lambda=2$. As we will see later, it is also useful to define $C_{1}=\mathcal{C}_{0}-D_{0}$ [i.e. we start the sum in Eq. (3) at $k=1$ rather than $k=0$]. We will also see that the contributions fall off rather rapidly with $k$, so that in practice only a limited number of terms need be considered.
@@ III Results
Figure 2: Classical models in $128\times 128$ lattices. Results are presented for Ising FM (a)-(b) and AFM (c)-(d) couplings, and for the XY model (e)-(f). Upper row: $\mathcal{C}_{0}$ (blue circles) and $\mathcal{C}_{1}$ (red squares) as a function of $T/J$. Solid lines correspond to the absolute value of the derivative. The extracted transition temperatures $T_{c}$ are consistent with the exact results in all cases. In panel (e) black markers and lines correspond to results obtained in $256\times 256$ lattices. Lower row: $D_{k}$ vs $k$ for different temperatures. The behavior of the dissimilarities for each model are signficantly different.
@@ III.1.1 Hamiltonians with $\mathcal{Z}_{2}$ symmetry
Finally, we notice that $\mathcal{D}_{0}$ in the FM Ising model closely tracks, within a prefactor, the temperature behavior of the entropy per spin $S$ (not shown), and $\mathcal{C}_{0}$ correlates with the Boltzmann entropy $S_{B}$, defined as [50]
$S_{B}=-\sum_{\sigma}p_{\sigma}\log_{2}(p_{\sigma}),$ (8)
where $p_{\sigma}=(1\pm m)/2$, and $m$ is the average magnetization of the system. That means $\mathcal{C}_{1}$ follows the trend in the total correlation density $\rho_{c}=S_{B}-S$ [50]. The Boltzmann entropy assumes the lack of internal correlations and corresponds to the isolated spin entropy. Since $\rho_{c}$ only vanishes when no internal correlations occur, i.e. when each spin is independent, it measures how constrained the spin distribution is, and is therefore maximized when the spins are strongly correlated [50].
@@ III.1.2 The Hamiltonian with $O(2)$ symmetry
We now turn our attention to the XY model. In Fig. 2(e) we show the structural complexity of the model as a function of temperature for two system sizes $128\times 128$ (solid markers) and $256\times 256$ (open black markers). The results are averaged over five different initial random seeds. We further apply a moving average with a five-point window over temperature fitted with a local third-order polynomial (the Savitzky-Golay filter) to reduce noise. Similarly to the AFM Ising model, $\mathcal{C}_{0}$ is more or less constant for all temperatures and for both system sizes, indicating the near zero net magnetization of the system. On the contrary, $\mathcal{C}_{1}$ increases as the temperature is lowered, indicating that the low-$T$ phase, where the existence of pairs of vortices with opposite vorticity is expected [47], is the most complex one. We find a strong finite-size dependence in $\mathcal{C}_{1}$, expected to persist in this model for certain properties even with systems an order of magnitude larger in size [51]. The derivative of $\mathcal{C}_{1}$ for the $128\times 128$ system exhibits broad and inconclusive features at $T/J\lesssim 3$. By increasing the system size to $256\times 256$, these features develop into a sharp upturn at the lowest temperatures we could access in our simulations due to inflating autocorrelation times. This feature is consistent with the location of the peak in the specific heat ($T/J\sim 1.04$) and the BKT transition temperature of the model ($T_{c}/J=0.894$[52, 51]. In order to access lower temperatures and larger system sizes, more sophisticated numerical algorithms are required, which are beyond the scope of this work [51, 53, 54].
@@ III.2.1 Spin-resolved densities $n_{\sigma}$
More interestingly, we find that $\mathcal{C}_{1}$ vs temperature or doping exhibits the same trends as the entropy per site $S/N$ in the model (see Fig. 7 for the doping dependence at different temperatures). We observe that below $T/t<1$, the structural complexity develops a peak around 20% doping. Although the peak in the entropy occurs at a slightly lower doping (see the inset of Fig. 7), the shape and trends are fairly similar. Additionally, we also observe good agreement with the experimental data at all temperatures considered, which further supports the validity of our results. Because of the correlation between $\mathcal{C}_{1}$ and $S/N$, this suggests the structural complexity of Fermi-Hubbard snapshots can be used as a quick and rapid proxy for the entropy of the system. This can be very useful in optical lattice experiments, where after loading the lattice, the calculation of the entropy is not straightforward [56].
It is worth mentioning that previous studies in one-dimensional systems: bit-strings [41] and the extended Hubbard model [43] have demonstrated that the structural complexity draws a close analog to the entanglement entropy since both capture the existence of nonlocal correlations in the models. In the present study we do not have access to the entanglement entropy and therefore only focus on the thermodynamic entropy. Although the entanglement entropy and the thermodynamic entropy are not equivalent to each other, they are associated with the development of correlations in the system and with the freezing of degrees of freedom, respectively, which generally occur together as we vary a tuning parameter. The establishment of connections between the structural complexity and the entanglement entropy in doped Fermi Hubbard models at finite-$T$ is an open question that requires further exploration.
Another important observation is the fact that at the temperatures considered, the region where the structural complexity is maximal corresponds to the strange metal region. The broad peak in $\mathcal{C}_{1}$ spanning $\delta\in[10\%,30\%]$ is consistent with the doping value $\delta_{c}\sim 30\%$ that corresponds to the crossover to the Fermi liquid phase [57]. Moreover, the trends in the structural complexity vs doping illustrates that features related to phases that emerge in this doping range at lower temperatures (such as fluctuating stripes [31, 33, 32]) are more complex than the AFM parent state at half-filling, and that this metallic region is also more complex than the conventional Fermi liquid found at larger dopings. As optical lattice experiments evolve and stripe physics becomes available, we expect that signatures of the presence and periodicity of stripe orderings should be detectable by analyzing the contributions of the dissimilarities at different scales and $\mathcal{C}_{1}$.
@@ III.2.2 Local moments and total density
Before quantum gas microscopy had access to images of both fermionic species of the same system at the same time, density snapshots of the FHM could not be obtained; only local moments could be detected separately. Nowadays, with spin-resolved imaging, local moment as well as total density images can be created [58, 28, 29]. In this section we study their structural complexities and dissimilarities.
@@
Fig. 11 presents $2^{\langle k\rangle}$ vs doping at different temperatures for the three different types of snapshots analyzed. At the highest temperature considered, $T/t=4$, $2^{\langle k\rangle}$ increases slowly with $\delta$ for the $n_{\sigma}$ and $n$ snapshots, but decreases for the $m_{z}$ snapshots. This trend persists down to all temperatures considered for the $m_{z}$ snapshots, becoming sharper as the temperature is lowered, but surprisingly exhibiting a crossing at $\delta\sim 18\%$, which points to a possible temperature-independent length scale in $m_{z}$ at this doping. This coincides with the region where structural complexity is maximal and corresponds to the strange metal region. Since this occurs at easily experimentally-accessible regimes, it is a phenomenon that can be explored immediately in the experiment.
$2^{\langle k\rangle}$ vs temperature for $n_{\sigma}$ and $n$ snapshots also displays interesting behaviors at the same doping. As seen in Fig. 11(a) for $n_{\sigma}$, the curves develop a plateau that starts around $\delta\sim 18\%$ and extends up to $\delta\sim 40\%$, while for $n$ in Fig. 11(b) the curves develop a broad peak around $\delta\sim 18\%$.
@@ Acknowledgements.
We are grateful to Waseem Bakr, Elmer Guardado-Sanchez, and Benjamin M. Spar for sharing their experimental snapshots for the Fermi-Hubbard model with us. We also thank Waseem Bakr, Martin Zwierlein for their useful comments on our manuscript. The authors are supported by the grant DE-SC-0022311, funded by the U.S. Department of Energy, Office of Science. EIGP would like to acknowledge useful discussions with Steven Johnston and Bejamin Cohen-Stead while visiting the University of Tennessee, Knoxville.